The following is a partially completed lower section of a departmental expense allocation spreadsheet for Brickland. It reports the total amounts of direct and indirect expenses for the four departments. Purchasing department expenses are allocated to the operating departments on the basis of purchase orders. Maintenance department expenses are allocated based on square footage. Compute the amount of Maintenance department expense to be allocated to Assembly.

Purchasing Maintenance Fabrication Assembly
Operating costs $ 37,000 $ 21,000 $ 101,000 $ 67,000
No. of purchase orders 14 6
Sq. ft. of space 3,550 2,450

Multiple Choice

  • $25,900.

  • $10,500.

  • $12,425.

  • $8,575.

  • $11,100.
Answer #1

Total square foot = 3,550 + 2,450 = 6,000

Maintenance department expense allocated to Assembly = ($21,000 / 6,000) X 2,450 = $8,575

4th Option.